When two given functions, f(x) = x


2 + 3 and g(x) = 2x − 5, are divided with each other, a


new third function is obtained. Find (f/g)(x) and identify its domain.


1
Expert's answer
2021-09-23T17:16:44-0400
"(f/g)(x)=\dfrac{x^2+5}{2x-5}"

"2x-5\not=0=>x\not=\dfrac{5}{2}"

"(f/g)(x)=\dfrac{x^2+5}{2x-5}"


"Domain: (-\infin, \dfrac{5}{2})\cup(\dfrac{5}{2}, \infin)"